OpenStudy (anonymous):

Does curve r1(s)=(s^4/2-3s^2/4,s^2,3s/4) and r2(t)=(t^2-t,2t,t^2+t) intersect? If they do, in how many points do they intersect?

OpenStudy (amistre64):

r1(s)= x=s^4/2-3s^2/4 y=0+s^2, z=0+3/4 s) eww, how would we write this out in parametric vorm?

OpenStudy (amistre64):

*form

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

i was thinking to equalize cooeficient and then i get s=0,t=0 and put it in r1, r2, i get r1=(0,0,0)r2=(0,0,0) so they do intersect in one point

OpenStudy (mr.math):

r_1(s)=r_2(t), iff s^4/2-3s^2/4=t^2-t s^2=2t 3s/4=t^2+t

OpenStudy (mr.math):

Use equation(2) to find t in terms of s, then substitute that in either equation 1 or 3 to find s.

OpenStudy (mr.math):

Make sure that the solution you get satisfy all three equations.

OpenStudy (mr.math):

The solutions are s=t=0 or s=1 and t=1/2. Plug that into r to find the coordinates of the points.